OpenStudy (anonymous):

What is the simplified form of the expression: 3 log 5 + 3 log 2 – 2 log 4 ?

OpenStudy (anonymous):

you should probably ask that in the math section.

OpenStudy (anonymous):

\[\log\left(\frac{125}{2} \right)\] Did you already get an answer in the math section? :)

OpenStudy (lanre):

log125+log8-log16=log117

OpenStudy (lanre):

sorry, log125+log8-log16=log125xlog8/log16=log125/log2

OpenStudy (anonymous):

???